Grade 11Physical Chemistry

A sample of sodium has a mass 46gm what is the mass of the same number of calcium atoms as sodium atoms present in given sample

anish vashishtha
10 Years agoGrade 11

4 Answers

Ashutosh Mohan Sharma
10 Years ago
atomic mass of sodium= 23 g/mol
moles of sodium atoms in 46 gm= 46/23 = 2 moles
as mentioned in question, no of Na atoms = no of atoms of Ca atoms.
therefore no of atoms of ca atoms= 2 moles
therefore mass of 2 moles of Ca atoms = 2 x atomic wt of Ca = 2 mol x 40 g/mol = 80 g
Therefore answer is 80g.
Shivam kumar
8 Years ago
Mass of one atom is 3.32×10×_23. Now weight of one mole is 3.32×10×_24.so number of nucleon present in 1 neutrons= 3.32×10×-23÷3.32×10×_24=10 neutrons in one atom.so no of neutrons present in 2 atom=10×2=20 neutrons
Anjali kumari
8 Years ago
As we know mole=given mass/molecular mass.So moles of sodium=2 moles.Now in 2 moles of calcium atom the mass is 80g.
Kartik garg
7 Years ago
Mass of sodium =46g
Molar mass of sodium=23g
No.  of moles of sodium =mass of sodium /molar mass of sodium 
                            =46/23=2moles of sodium 
Same no of calcium atom as sodium atom
Molar mass of calcium=40g
                    No. of moles=mass/molar mass
                      Massof calcium =2*40=80g